Three times the complement of an angle less the supplement of an angle is 10 degrees. Find the supplement.
Answer by edjones(7569)

let x=the angle
supplement=180-x
complement=90-x
3(90-x)-(180-x)=10
270-3x-180+x=10
90-2x=10
subtract 90 from each side: -2x=-80
divide -2 into each side: x=40
Check:
3*50-140=10
150-140=10
Ed